OpenStudy (anonymous):

Write a proportion for the following problem. The tax on a new car costing $14,400 is $784. At the same tax rate, what is the tax on a vehicle costing $20,000?

OpenStudy (anonymous):

ur not right because ur answer isnt one of my options

OpenStudy (anonymous):

84/14,400 = x/20,000 20,000/14,400=784/x 784/14,400 = 20,000/x

OpenStudy (nurali):

784/14,400 = x/20,000

OpenStudy (anonymous):

@nurali is right 784/14,400 = x/20,000 So the answer is x = 1089